OpenStudy (anonymous):

the odds against lounging larry winning the horse race are 5:8. determine the probability that lounging larry wins the horse race.

OpenStudy (heisenberg):

The probability that something happens + the probability that something doesn't happen = 1

OpenStudy (heisenberg):

if the probability that larry doesn't win the horse race is 5/8, what is the probability he wins?

OpenStudy (anonymous):

13/5

OpenStudy (heisenberg):

that doesn't seem right. it would equals 1 - 5/8

OpenStudy (heisenberg):

1 could also be written as \(\frac{8}{8}\)

OpenStudy (anonymous):

is it 8/13

OpenStudy (heisenberg):

no, not that either. do you know how to subtract fractions? can you solve: \[\frac{8}{8} - \frac{5}{8} \]

OpenStudy (heisenberg):

...wait i may be wrong :/ heh

OpenStudy (heisenberg):

if the odds that larry loses are 5:8 then that means the probability he loses is 5/13, I believe. which would make 8/13 correct (what you said initially)

OpenStudy (heisenberg):

because odds means x:y where x is one outcome and y is another. probably is x/y where x is one outcome and y is all possible outcomes.

OpenStudy (anonymous):

ok got it thanks alot